he makes  machine. Here is the relationship between the number of workers and he ’s output during a given day:

 

Workers

Output

Marginal Product

Total Cost

Average Total Cost

Marginal Cost

0

0

--

 

--

--

1

20

 

 

 

 

2

50

 

 

 

 

3

90

 

 

 

 

4

120

 

 

 

 

5

140

 

 

 

 

6

150

 

 

 

 

7

155

 

 

 

 

 

1) Fill in the column of marginal product. What pattern do you see? How might you explain it?

2) A worker costs $100 a day, and the firm has fixed costs of $200. Use this information to fill in the column for TC (Total Cost).

3) Fill in the column for ATC (Average Total Cost). What pattern do you see? (cf. ATC=TC/Q)

4) Now fill in the column for marginal cost. What pattern do you see? (cf. MC=△TC/△Q)

 


1
Expert's answer
2021-06-07T10:34:50-0400


From the above table;

1)

"MP=\frac{\Delta output}{\Delta workers}"

The marginal cost increases with increase in the number of workers reaching maximum (40) when the number of workers is 3 and then starts to drop.

This can be explained using the law of diminishing returns which states that; if more variable input units are used along with a certain amount of fixed inputs, the overall output might grow at a faster rate initially, then at a steady rate, but ultimately, it will grow at a declining rate.


2.)

Let no of workers =n

"TFC=200\\

AVC=100\\

TVC=100n\\

TC=TVC+TFC"

"TC=100n+200"



3.)

"ATC=\frac{TC}{Q}"

ATC reduces with the increase in output reaching minimum (5) at a point where number of outputs is 120 and 140 then it starts increasing again.


4.)

Marginal cost decreases with increase in output reaching minimum (2.5) when output is 90 and then starts increasing again.